# Nearly-polynomial inverse theorem for the $U^{d}$ norm in degree $d+1$

### 17 Mar 2026 v1

@@ 1 Introduction
Let $\mathbb{F}=\mathbb{F}_{p}$ be any prime finite field. The discrete derivative of a function $f\colon\mathbb{F}^{n}\to\mathbb{F}$ in direction $v\in\mathbb{F}^{n}$ at the point $x\in\mathbb{F}^{n}$ is $\Delta_{v}f(x)=f(x+v)-f(x)$. More generally, the (discrete) derivative of order $d$, in directions $\mathbf{v}=(v^{(1)},\ldots,v^{(d)})\in(\mathbb{F}^{n})^{d}$ at the point $x\in\mathbb{F}^{n}$, is
@@
A function $f\colon\mathbb{F}^{n}\to\mathbb{F}$ is a polynomial of degree below $d$, denoted $f\in\operatorname{Poly}_{<d}$, if and only if $\Delta^{d}f=0$ (that is, $\Delta^{d}_{\mathbf{v}}f(x)=0$ for every $(x,\mathbf{v})\in(\mathbb{F}^{n})^{d+1}$).
An important (qualitative) result about polynomials over $\mathbb{F}$, of characteristic $\operatorname{ch}(\mathbb{F})>d$, is that the derivative characterization is robust: $f$ is correlated with $\operatorname{Poly}_{<d}$ if and only if $\Delta^{d}f$ is biased. This lies at the core of the theory of Gowers norms [Gow01, Tao12, HHL19], and is usually formulated using exponential sums: [^1] Which measure the distance of a distribution to uniformity; see e.g. Claim 33 in [BV10]. for any function $f\colon\mathbb{F}^{m}\to\mathbb{F}$, one denotes
@@
where $e(y):=e^{2\pi iy/p}\colon\mathbb{F}\to\mathbb{C}$ and $\operatorname*{\mathbb{E}}_{x\in A}:=\frac{1}{|A|}\sum_{x\in A}$. We note that $\operatorname{bias}(f)\in[0,1]$, [^2] For example, if $\deg(f)<d$ then $\operatorname{bias}(\Delta^{d}f)=1$, while if $f$ is uniformly random then $\operatorname{bias}(\Delta^{d}f)=o(1)$. and that, explicitly, $\operatorname{bias}(\Delta^{d}f):=\operatorname*{\mathbb{E}}_{x\in\mathbb{F}^{n},\mathbf{v}\in(\mathbb{F}^{n})^{d}}e(\Delta^{d}_{\mathbf{v}}f(x))$. The forward direction of this characterization—correlation implies bias—is not hard to prove, even quantitatively (e.g., using the Cauchy-Schwarz-Gowers inequality):
@@
The inverse direction is much harder. For $d=2$, the question falls in the domain of Fourier analysis, as $\operatorname{bias}(\Delta^{2}f)=\sum_{\zeta}|\hat{f}(\zeta)|^{4}$, and Parseval’s identity implies $\operatorname{cor}_{<2}(f)\geq\operatorname{bias}(\Delta^{2}f)^{1/2}$. However, standard Fourier analysis is insufficient for general $d$, and the inverse direction is only known to hold with an extremely weak bound [GM20]. The inverse direction is often stated in terms of the Gowers (uniformity) norms of a complex-valued function. For $g\colon\mathbb{F}^{n}\to\mathbb{C}$, the $U^{d}$-norm is $\lVert{g}\rVert_{U^{d}}=(\operatorname*{\mathbb{E}}_{(x,\mathbf{v})\in(\mathbb{F}^{n})^{d+1}}{\Delta^{*}}_{v^{(d)}}\cdots{\Delta^{*}}_{v^{(1)}}g(x))^{1/2^{d}}$, with $\Delta^{*}_{v}g(x)=g(x+v)\overline{g(x)}$, and the weak Gowers norm is $\lVert{g}\rVert_{u^{d}}=\max_{\deg(P)<d}\big|\operatorname*{\mathbb{E}}_{x\in\mathbb{F}^{n}}g(x)e(-P(x))\big|$.
Theorem 1.1 (Gowers inverse over finite fields [GT09, BTZ10, TZ10, TZ12]).
@@
We note that if the field has small characteristic $\operatorname{ch}(\mathbb{F})<d-1$, the maximum in Theorem 1.1, over $P$ with $\deg(P)<d$, must also range over so-called non-classical polynomials of degree below $d$ ([BS+22, GT09, LMS11]), which are functions $\mathbb{F}^{n}\to\mathbb{R}/\mathbb{Z}$ that involve division by powers of $\operatorname{ch}(\mathbb{F})$. In this paper we restrict ourselves to the high characteristic case (and so, to classical polynomials).
The Polynomial Gowers Inverse conjecture posits that, in fact, the relationship between bias and correlation is about as tight as it could be; namely, that $\Phi(x)$ in Theorem 1.1 can be as large as a power of $x$ for any fixed $d$: $\Phi(x)\geq x^{O_{d}(1)}$ (this matches the easy direction $\Phi(x)\leq x^{2^{d}}$ in (1) up to the constant). It is perhaps the central conjecture of higher-order Fourier analysis, over finite fields and over the integers (e.g. [LSS24]), with applications in countless fields (see, e.g., [HHL19, Tao12, TZ10] and the references within).
@@ 1.1 Gowers inverse in low degree
In this paper we obtain an almost-polynomial bound for $\operatorname{GI}_{d}(d+1)$, the next open case. With a slight abuse of notation, we write $x^{\tilde{O}_{d}(1)}$ for $\exp(-O_{d}(\log(1/x)^{1+o(1)}))$. (This is even stronger than a quasi-polynomial bound, which has a constant $O_{d}(1)$ at the top instead of 1+o(1)1+o(1).) [^4] More precisely, we get $\log(1/x)\cdot\log(\log(1/x)+1)$ in the exponent. (In fact, we even get a bound that improves as $|\mathbb{F}|$ increases, of the rough shape $x^{O_{d}(\log_{|\mathbb{F}|}(\log_{|\mathbb{F}|}(1/x)))}$, which for $|\mathbb{F}|\gg\log(1/x)$ is truly polynomial $x^{O_{d}(1)}$; see [MZ22].)
@@
The main difficulty in the proof of Theorem 1 stems from the non-homogeneity of $f$. Standard proofs of the degree-$d$ case proceed by finding a decomposition of $f\in\operatorname{Poly}_{d}$ into lower-degree polynomials of length bounded in terms of $\operatorname{bias}(\Delta^{d}f)$, which is then used to obtain a significant correlation with a lower-degree polynomial. However, when $\deg(f)=d+1$, one obtains in this way only a decomposition of the top-degree part of $f$, and indeed, it is entirely possible for the degree-$d$ part $f_{d}$ of $f$ to not have any short decompositions despite having large $\operatorname{bias}(\Delta^{d}f)$ (e.g., $f=(x_{1}+1)h$ for a random $h$, see Remark 6.2). Furthermore, even if one were to successfully obtain a short decomposition of $f$ in terms of polynomials of degrees below $\deg(f)$, it is not clear how to use this to obtain a correlation with a polynomial of degree as small as $d-1=\deg(f)-2$; especially given that, unlike in the degree-$d$ case, one cannot guarantee a short decomposition with homogeneous factors. [^5] Concretely, the degree-$1$ factors in a decomposition of $f$ might not have a common zero (as they need not be homogeneous; e.g., $x_{1}$ and $x_{1}+1$), and so there might not even be a large subspace, or affine subspace, where the degree-$1$ factors can be discarded (which would have meant that all factors have degree at most $\deg(f)-2$).
In the homogeneous setting, we can do more. Namely, we obtain almost-polynomial bounds for homogeneous polynomials of degree up to $2d$.
@@
$\operatorname{cor}_{<d}(f)\geq\operatorname{bias}(\Delta^{d}f)^{\tilde{O}_{d}(1)}$
The proof of Theorem 2 relies on a new, general correlation lemma. Unlike standard correlation lemmas in the literature (e.g. [Jan19], proof of Theorem 1.6, or [GT09, CM23] and many more) which require a full decomposition of the given polynomial, ours requires a more relaxed condition, closely related to membership in an appropriate ideal.
@@ 2.1 Preliminaries
Henceforth, $\mathbb{F}$ is any finite field. We write $\operatorname{\mathfrak{bias}}$ (in Fraktur letters) for the bias as a complex number: for any function $f\colon\mathbb{F}^{n}\to\mathbb{F}$ over a finite field $\mathbb{F}$, denote
@@
A form is a homogeneous polynomial. We denote by $\operatorname{Form}_{k}(\mathbb{F})$ the set of forms of degree $k$, over the field $\mathbb{F}$, in any (finite) number of variables, or simply $\operatorname{Form}_{k}$ when $\mathbb{F}$ is immaterial. Similarly, we denote by $\operatorname{Poly}_{k}(\mathbb{F})$ the set of polynomials of degree at most $k$ over $\mathbb{F}$ is any number of variables. Note that any $f\in\operatorname{Poly}_{k}$ is the sum of forms, namely, the homogeneous components of $f$ of each degree: $f=\sum_{i=0}^{k}f_{i}$ with $f_{i}\in\operatorname{Form}_{i}$. We also denote $f_{<d}:=\sum_{i=0}^{d-1}f_{i}$.
@@
for some forms $g_{i}$ and $h_{i}$ of degree below $d$. A form is multilinear if it is linear in each of its variable vectors; more concretely, a multilinear form of degree $d$, or a $d$-linear form for short, is of the form $T=T(x^{(1)},\ldots,x^{(d)}):(\mathbb{F}^{n})^{d}\to\mathbb{F}$ and is linear in each of $x^{(1)},\ldots,x^{(d)}$. Equivalently, any monomial in the support of $T$ has one variable from each variable vector. The partition rank [Nas20] of a $d$-linear form $T$, denoted by $\operatorname{PR}(T)$, is the minimal $r$ for which $T$ is the sum of reducible $d$-linear forms:
@@
for some $R_{i}=R_{i}(x_{I_{i}})$ and $Q_{i}=Q_{i}(x_{I_{i}^{c}})$ with $\emptyset\subsetneq I_{i}\subsetneq[d]$, where for $x\in(\mathbb{F}^{n})^{d}$ we write $x_{I}$ for the restriction of $x$ to the coordinates corresponding to $I\subseteq[d]$. A related notion to the partition rank is that of the analytic rank [GW11] over a finite field $\mathbb{F}$, denoted by $\operatorname{AR}(T)$:
@@
We note that for a multilinear $T$, we have $\operatorname{bias}(T)=\mathbb{P}(T=0)-\mathbb{P}(T=c)$ for any nonzero $c\in\mathbb{F}$.
The polarization of a form $g\in\operatorname{Form}_{k}$, which we denote by $\overline{g}$, is the unique (degree-$k$) multilinear form satisfying $\overline{g}(x,\ldots,x)=k!g(x)$. (Note that if $g\colon\mathbb{F}^{n}\to\mathbb{F}$ then $\overline{g}\colon(\mathbb{F}^{n})^{k}\to\mathbb{F}$.) Equivalently, $\overline{g}$ can be obtained as $\overline{g}=\Delta^{k}g$, the discrete derivative of order $k$; that is, $\overline{g}(\mathbf{x})=\Delta^{k}_{\mathbf{x}}g$ where $\mathbf{x}=(x^{(1)},\ldots,x^{(k)})$. (See Fact 5.7 below.) Note that $\Delta^{k}_{\mathbf{x}}g$ depends only on $\mathbf{x}$, being independent of the point at which we take the discrete derivative (as $k=\deg(g)$). For example, for $g(x)=x_{1}x_{2}$, we have $\overline{g}(x,y)=x_{1}y_{2}+y_{1}x_{2}$, and for $g(x)=x_{1}^{2}$ we have $\overline{g}(x,y)=2x_{1}y_{1}$.
@@
A fundamental part of our argument relies on the best bounds known for the rank of biased homogeneous polynomials (see, e.g., Corollary 8.5 in [MZ22] [^6] One can naively extend the definition of $\operatorname{rk}$ to non-homogeneous polynomials $f\in\operatorname{Poly}_{k}$ by defining $\operatorname{rk}(f):=\operatorname{rk}(f_{k})$, but this is not useful in this paper as this does not give a decomposition of $f$, only of $f_{k}$.). We write $\tilde{O}_{d}(x):=O_{d}(x^{1+o(1)})$. [^7] More precisely, we can get $\tilde{O}_{d}(x)=O_{d}(x\log(x+1))$. (In fact, we can even get $\tilde{O}_{d}(x)=O_{d}(x(\log_{|\mathbb{F}|}(x+1)+1))$, which in particular is linear for mildly large $\mathbb{F}$).
Theorem 2.1 ([MZ22]).
For every form $g\in\operatorname{Form}_{d}$, over any finite field $\mathbb{F}$ with $\operatorname{ch}(\mathbb{F})>d$,
$\operatorname{rk}(g)\leq\tilde{O}_{d}(\operatorname{AR}(g)).$
Moreover, for every $d$-linear form $T$,
$\operatorname{PR}(T)=\tilde{O}_{d}(\operatorname{AR}(T))$

### 30 Apr 2026 v2

@@ Abstract
Our methods may be of independent interest, and include a refined notion of polynomial decomposition that captures correlation with polynomials of lower degree than classical notions do, and a new correlation lemma that improves upon similar lemmas in the literature. Additionally, we illustrate the usefulness of the new correlation lemma by using it to give an alternative proof for the aforementioned result of Milićević and Rand¯elović.
@@ 1 Introduction
Let $\mathbb{F}$ be a finite field, for simplicity prime $\mathbb{F}=\mathbb{F}_{p}$. The discrete derivative of a function $f\colon\mathbb{F}^{n}\to\mathbb{F}$ in direction $v\in\mathbb{F}^{n}$ at the point $x\in\mathbb{F}^{n}$ is $\Delta_{v}f(x)=f(x+v)-f(x)$. More generally, the (discrete) derivative of order $d$, in directions $\mathbf{v}=(v^{(1)},\ldots,v^{(d)})\in(\mathbb{F}^{n})^{d}$ at $x\in\mathbb{F}^{n}$, is
@@
We write $\Delta^{d}f$ for the map $(x,\mathbf{v})\mapsto\Delta^{d}_{\mathbf{v}}f(x)$. A function $f\colon\mathbb{F}^{n}\to\mathbb{F}$ can be written [^1] Or: the unique reduced polynomial (variable degrees below $|\mathbb{F}|$) agreeing with $f$ has (total) degree below $d$. as a polynomial of degree below $d$ if and only if $\Delta^{d}f=0$ (i.e., $\Delta^{d}_{\mathbf{v}}f(x)=0$ for every $(x,\mathbf{v})\in(\mathbb{F}^{n})^{d+1}$).
An important (qualitative) result about polynomials over $\mathbb{F}$, of characteristic $\operatorname{ch}(\mathbb{F})>d$, is that the derivative characterization is robust: $f$ is correlated with some polynomial of degree below $d$ if and only if $\Delta^{d}f$ is biased. This lies at the core of the theory of Gowers norms [Gow01, Tao12, HHL19], and is usually formulated using exponential sums: [^2] Which are closely related to the distance of a distribution from uniformity; see e.g. Claim 33 in [BV10]. for any function $f\colon\mathbb{F}^{m}\to\mathbb{F}$, one denotes
@@
where $e(y):=e^{2\pi iy/p}\colon\mathbb{F}\to\mathbb{C}$ and $\operatorname*{\mathbb{E}}_{x\in A}:=\frac{1}{|A|}\sum_{x\in A}$. We note that $\operatorname{bias}(f)\in[0,1]$, [^3] For example, if $\deg(f)<d$ then $\operatorname{bias}(\Delta^{d}f)=1$, while if $f$ is random then $\operatorname{bias}(\Delta^{d}f)=o(1)$ as $n\to\infty$. and that, explicitly, $\operatorname{bias}(\Delta^{d}f)=\operatorname*{\mathbb{E}}_{x\in\mathbb{F}^{n},\mathbf{v}\in(\mathbb{F}^{n})^{d}}e(\Delta^{d}_{\mathbf{v}}f(x))$. [^4] The absolute value is not needed for $\Delta^{d}f$. The forward direction of this characterization—correlation implies bias—is not hard to prove, even quantitatively (e.g., using [^5] Indeed, if $\deg(P)<d$ then $\operatorname{bias}(\Delta^{d}f)=\operatorname{bias}(\Delta^{d}(f-P))\geq|\operatorname{bias}(f-P)|^{2^{d}}=\operatorname{cor}(f,P)^{2^{d}}$. the Cauchy-Schwarz-Gowers inequality or the monotonicity of Gowers norms [Tao12]):
@@
The inverse direction is much harder. For $d=2$, the question falls in the domain of Fourier analysis, and it follows using Parseval’s identity that $\operatorname{cor}_{<2}(f)\geq\operatorname{bias}(\Delta^{2}f)^{1/2}$. [^6] Indeed, the magnitudes of the Fourier coefficients of $g=e(f)$ are $\operatorname{cor}(f,\ell)$, for any linear form $\ell$ (or character), and $\operatorname{bias}(\Delta^{2}f)=\sum_{\gamma}|\hat{g}(\gamma)|^{4}=\sum_{\ell}\operatorname{cor}(f,\ell)^{4}\leq\operatorname{cor}_{<2}(f)^{2}\sum_{\ell}\operatorname{cor}(f,\ell)^{2}=\operatorname{cor}_{<2}(f)^{2}$. For general $d$, however, the inverse question falls in the domain of higher-order Fourier analysis, and is only known to hold with a very weak bound [GM20] (stronger bounds are known in the integer setting [LSS24]). The inverse question is often stated in terms of the Gowers (uniformity) norms of a complex-valued function. For $g\colon\mathbb{F}^{n}\to\mathbb{C}$, the $U^{d}$-norm is $\lVert{g}\rVert_{U^{d}}=(\operatorname*{\mathbb{E}}_{(x,\mathbf{v})\in(\mathbb{F}^{n})^{d+1}}{\Delta^{*}}_{v^{(d)}}\cdots{\Delta^{*}}_{v^{(1)}}g(x))^{1/2^{d}}$, with $\Delta^{*}_{v}g(x)=g(x+v)\overline{g(x)}$, and the weak Gowers norm is $\lVert{g}\rVert_{u^{d}}=\max_{\deg(P)<d}\big|\operatorname*{\mathbb{E}}_{x\in\mathbb{F}^{n}}g(x)e(-P(x))\big|$.
Theorem 1.1 (Gowers inverse over finite fields [BTZ10, TZ10, TZ12]).
@@
We note that if the field has small characteristic $\operatorname{ch}(\mathbb{F})<d-1$, the maximum in Theorem 1.1, over $P$ with $\deg(P)<d$, must also range over so-called non-classical polynomials of degree below $d$ (see [TZ12] and [BS+22, GT09, LMS11]), which are functions $\mathbb{F}^{n}\to\mathbb{R}/\mathbb{Z}$ that involve division by powers of $\operatorname{ch}(\mathbb{F})$. In this paper we restrict ourselves to the high characteristic case (and so, to classical polynomials).
The Gowers inverse question goes back to foundational work of Gowers [Gow01], which led to the development of higher-order Fourier analysis, and seminal works of Green and Tao on additive patterns in the primes [GT10] and the $U^{3}$-norm [GT08], and of Samorodnitsky [Sam07] on polynomiality testing. It has found applications in a large number of areas (see, e.g., [BV10, HHL19, Pel24, Tao12, TZ10, Wol15] and the references within). However, the lack of strong bounds for Theorem 1.1 implies the same for the applications.
The Polynomial Gowers Inverse conjecture [GT08] posits that, in fact, the relationship between bias and correlation is about as tight as it could be. Namely, the conjecture is that $\Phi(x)$ in Theorem 1.1 can be as large as a power of $x$ for any fixed $d$: $\Phi(x)\geq x^{O_{d}(1)}$ (this matches the easy direction $\Phi(x)\leq x^{1/2^{d}}$ in (1) up to the constant). It is perhaps the central conjecture of higher-order Fourier analysis.
@@ 1.1 Gowers inverse in low degree
In this paper we obtain an almost-polynomial bound for $\operatorname{GI}_{d}(d+1)$, the next open case. With a slight abuse of notation, given $x\in(0,1]$ we write $x^{\tilde{O}_{d}(1)}$ for $\exp(-O_{d}(\log(1/x)^{1+o(1)}))$. [^9] For completeness, set $0^{\tilde{O}_{d}(1)}=0$. (This is even stronger than a quasi-polynomial bound, which has a constant $O_{d}(1)$ at the top instead of 1+o(1)1+o(1).) [^10] More precisely, we get $\log(1/x)\cdot\log(\log(1/x)+1)$ in the exponent. (We even get a bound that improves as $|\mathbb{F}|$ increases, of the shape $x^{O_{d}(\log_{|\mathbb{F}|}(\log_{|\mathbb{F}|}(1/x)))}$, which for $|\mathbb{F}|\gg\log(1/x)$ is truly polynomial $x^{O_{d}(1)}$; see [MZ22].)
@@
A main difficulty in the proof of Theorem 1 stems from the non-homogeneity of $f$. Standard proofs of the degree-$d$ case utilize the fact that if $\deg(f)\leq d$, then the $d$-times derivative $\Delta^{d}f$ is a multilinear form, a class of polynomials for which we have good understanding of the interplay between structure and randomness. Concretely, we know how to convert high $\operatorname{bias}(\Delta^{d}f)$ into a decomposition of $f_{d}$—the top-degree homogeneous component of ff—in terms of few lower-degree polynomials. A standard pigeonhole argument is then used to obtain a significant correlation for $f_{d}$ with a polynomial $P$ of lower degree, $\deg(P)\leq d-1$, and thus for $f$ itself by absorbing its lower-degree components into $P$.
However, when $\deg(f)=d+1$ and $f$ is not homogeneous, one obtains from the multilinear argument (via $\Delta^{d+1}f$, which is also biased) only a decomposition of the top-degree component $f_{d+1}$ of $f$, so absorbing the rest of ff—which is of degree dd—into a correlating polynomial $P$ with $\deg(P)\leq d-1$ is no longer an option. In fact, it is entirely possible for the degree-$d$ component $f_{d}$ of $f$ to not have any short decompositions, even if $\operatorname{bias}(\Delta^{d}f)$ is large (e.g., $f=(x_{1}+1)h$ for a random $h$, see Remark 6.2). Furthermore, even if one were to successfully obtain a short decomposition of $f$ itself in terms of polynomials of degrees below $\deg(f)$, it is not clear how to use this to obtain a correlation with a polynomial of degree as small as $d-1=\deg(f)-2$. [^11] Moreover, unlike in the degree-$d$ case, one cannot guarantee a short decomposition with homogeneous factors. In particular, the degree-$1$ factors in a decomposition of $f$ might not have a common zero (as they need not be homogeneous; e.g., $x_{1}$ and $x_{1}+1$), and so there might not even be a large (affine) subspace where the degree-$1$ factors can be discarded (which would have at least meant that all factors have degree at most $\deg(f)-2$ there).
In the homogeneous setting, we can do more. Namely, we obtain almost-polynomial bounds for homogeneous polynomials of degree smaller than $2d$.
@@
The proof of Theorem 2 relies on a new, general correlation lemma. Standard correlation lemmas/arguments in the literature (e.g. [Jan20], proof of Theorem 1.6; or [GT09], beginning of the proof of Theorem 1.4; or [KL08], proof of Theorem 3; or [TZ12], beginning of the proof of Theorem 1.10 from Theorem 1.18; and many more) use a simple pigeonhole-type argument on Fourier coefficients to obtain significant correlation with a lower degree polynomial, provided we have a full decomposition of the given polynomial. Our new correlation lemma, on the other hand, requires a more relaxed condition, closely related to membership in an appropriate ideal.
@@ 2.1 Preliminaries
From now on $\mathbb{F}$ is an arbitrary finite field. We write $\operatorname{\mathfrak{bias}}$ (in Fraktur letters) for the bias as a complex number: for any function $f\colon\mathbb{F}^{n}\to\mathbb{F}$ over a finite field $\mathbb{F}$, denote
@@
We denote by $\operatorname{Poly}_{k}(\mathbb{F})$ the set of (formal) polynomials of degree at most $k$, over the field $\mathbb{F}$, in any (finite) number of variables $x_{1},\ldots,x_{n}$. [^12] When combining polynomials, they are understood to be defined on the same variable set (adding dummy variables if needed). A form is a homogeneous polynomial. We denote by $\operatorname{Form}_{k}(\mathbb{F})$ the set of forms of degree $k$ ($\geq 0$) over $\mathbb{F}$ in any number of variables. We simply write $\operatorname{Form}_{k}$ when $\mathbb{F}$ is immaterial, or $\operatorname{Form}$ when $k$ is immaterial as well. Note that any $f\in\operatorname{Poly}_{k}$ is the sum of forms, namely, the homogeneous components of $f$ of each degree: $f=\sum_{i=0}^{k}f_{i}$ with $f_{i}\in\operatorname{Form}_{i}$. We also denote $f_{<d}:=\sum_{i=0}^{d-1}f_{i}$.
@@
for some forms $g_{i}$ and $h_{i}$ of (positive) degree smaller than $\deg(f)$. A form is multilinear if it is linear in each of its variable vectors; more concretely, a multilinear form of degree $d$ ($\geq 1$), or a $d$-linear form for short, is of the form $T=T(x^{(1)},\ldots,x^{(d)}):(\mathbb{F}^{n})^{d}\to\mathbb{F}$ and is linear in each of $x^{(1)},\ldots,x^{(d)}$. Equivalently, any monomial in the support of $T$ has one variable from each of $x^{(1)},\ldots,x^{(d)}$. The partition rank [Nas20] of a $d$-linear form $T$, denoted by $\operatorname{PR}(T)$, is the minimal $r$ for which $T$ is the sum of reducible $d$-linear forms:
@@
for some multilinear forms $R_{i}=R_{i}(x_{I_{i}})$ and $Q_{i}=Q_{i}(x_{I_{i}^{c}})$ with $\emptyset\subsetneq I_{i}\subsetneq[d]$, where for $x\in(\mathbb{F}^{n})^{d}$ we write $x_{I}$ for the restriction of $x$ to the coordinates corresponding to $I\subseteq[d]$. A related notion to the partition rank is that of the analytic rank [GW11] over a finite field $\mathbb{F}$, denoted by $\operatorname{AR}(T)$:
@@
We note that for $T$ multilinear, $\operatorname{bias}(T)$ also has an elementary definition: $\operatorname{bias}(T)=\mathbb{P}(T=0)-\mathbb{P}(T=c)$ ($\geq 0$), for any choice of nonzero $c\in\mathbb{F}$. [^13] Indeed, multilinearity implies that $\mathbb{P}(T=t)$ is the same for every $t\in\mathbb{F}^{\times}$, so $\operatorname{\mathfrak{bias}}(T)=\sum_{t\in\mathbb{F}}\mathbb{P}(T=t)\chi(t)=\mathbb{P}(T=0)+\mathbb{P}(T=c)(\sum_{t\in\mathbb{F}}\chi(t)-\chi(0))=\mathbb{P}(T=0)-\mathbb{P}(T=c)$. Moreover, $\operatorname{AR}(T)\leq\operatorname{PR}(T)$ (e.g., [Lov19]).
A fundamental part of our argument relies on the best bound known for the partition rank of biased multilinear forms. [^14] For completeness, note that if $d=1$ then $\operatorname{AR}(T)=\operatorname{PR}(T)=\infty$. We write $\tilde{O}_{d}(x):=O_{d}(x^{1+o(1)})$. [^15] More precisely, we can get $\tilde{O}_{d}(x)=O_{d}(x\log(x+1))$. (In fact, we can even get $\tilde{O}_{d}(x)=O_{d}(x(\log_{|\mathbb{F}|}(x+1)+1))$, which in particular is linear for mildly large $\mathbb{F}$).
Theorem 2.1 ([MZ22]).
For every $d$-linear form $T$ over any finite field,
$\operatorname{PR}(T)=\tilde{O}_{d}(\operatorname{AR}(T)).$
The polarization of a form $g\in\operatorname{Form}_{k}$, which we denote by $\overline{g}$, is the $k$-linear form $\overline{g}=\Delta^{k}g$, given by the discrete derivative of order $k$; that is, $\overline{g}(\mathbf{x})=\Delta^{k}_{\mathbf{x}}g$ where $\mathbf{x}=(x^{(1)},\ldots,x^{(k)})$. (Note that if $g\colon\mathbb{F}^{n}\to\mathbb{F}$ then $\overline{g}\colon(\mathbb{F}^{n})^{k}\to\mathbb{F}$; indeed, $\Delta^{k}_{\mathbf{x}}g$ depends only on $\mathbf{x}$, being independent of the point at which we take the discrete derivative, as $k=\deg(g)$.) If $k<\operatorname{ch}(\mathbb{F})$, $g$ can be recovered from its polarization: $g(x)=\overline{g}(x,\ldots,x)/k!$ (and $\overline{g}$ is the unique symmetric $k$-linear form satisfying this); see Fact 5.7 below. For example, for $g(x)=x_{1}x_{2}$, we have $\overline{g}(x,y)=x_{1}y_{2}+y_{1}x_{2}$, and for $g(x)=x_{1}^{2}$ we have $\overline{g}(x,y)=2x_{1}y_{1}$ (so $\overline{g}(x,x)$ fails to recover $g$ only when $\operatorname{ch}(\mathbb{F})=2$).